OpenStudy (anonymous):

Find the center and the radius of the circle x^2+y^2-12x+6y=19

OpenStudy (amistre64):

wolfram it :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x^2+y^2-12x+6y+36-36+9-9=19 x^2-12x+36 +y^2+6y+9=19+45 (x-6)^2 +(y+3)^2 =64=8^2 Centre = (6,-3) Radius =8

OpenStudy (anonymous):

Oh thanks! :D

OpenStudy (anonymous):

Welcome!

OpenStudy (amistre64):